# The Open mulTiwavelength Transient Event Repository (OTTER): Infrastructure Release and Tidal Disruption Event Catalog

### v1

@@ Appendix A Metadata and References for each TDE
ASASSN-15lh 22:02:15.49 -61:39:33.6 0.2326 2015-06-14 3.3 41 , 51 , 59 , 74 , 84 , 86 , 98 , 100 , 120 , 346 , 379
ASASSN-15oi 20:39:08.93 -30:45:22.6 0.0484 2015-08-14 3.3 5 , 60 , 74 , 83 , 98 , 100 , 105 , 198 , 201 , 244 , 346 , 371 , 376 , 379 , 387
OGLE16aaa 01:07:20.76144 -64:16:20.4456 0.1655 2015-12-17 3.3 5 , 74 , 98 , 100 , 107 , 108 , 346 , 379
Gaia 16aax 14:34:18 +12:49:37 0.248 2016-01-26 3.3 107 , 196 , 346
Gaia 16ajq 14:53:02 +42:21:28 0.28 2016-03-29 3.3 88 , 89 , 107 , 130 , 346
Gaia 16aka 14:48:08 +37:13:59 0.31 2016-03-31 3.3 88 , 107 , 346
iPTF16axa 17:03:34.36 +30:35:36.8 0.108 2016-05-29 3.3 5 , 74 , 98 , 101 , 111 , 346 , 373 , 379
2016fnl 00:29:57.04 +32:53:37.5 0.0163 2016-08-26 3.3 5 , 74 , 75 , 87 , 91 , 98 , 100 , 104 , 111 , 239 , 346 , 373 , 376 , 379
AT2016ezh 01:58:05 -00:52:22 0.0804 2016-08-26 3.3 5 , 51 , 74 , 90 , 98 , 103 , 114 , 202 , 236 , 346 , 373
OGLE17aaj 01:56:24.93 -71:04:15.7 0.116 2017-01-01 3.1 74 , 94 , 97 , 98 , 149 , 346
2017bcc 11:31:52.9875888 +29:59:44.8267578 0.133 2017-02-18 3.3 74 , 97 , 98 , 117 , 118 , 346 , 382
2019pqt 18:08:07.549 +13:15:05.4 2017-03-15 0.0 95 , 96 , 98 , 183 , 242 , 346
2017eqx 22:26:48.37042128 +17:08:52.39825116 0.1089 2017-06-01 3.3 74 , 98 , 119 , 160 , 202 , 346 , 373 , 379
2018gn 01:46:42.4356 +32:30:29.4048 0.037500 2018-01-14 3.3 133 , 138 , 368
XMMSL2 J140446.9-251135 14:04:47 -25:11:35 0.043 2018-02-15 1.0 122 , 346
2018zr 07:56:54.54888 +34:15:43.6284 0.071 2018-03-06 3.3 74 , 98 , 121 , 139 , 153 , 155 , 156 , 202 , 205 , 340 , 346 , 371 , 373 , 376 , 379
2018bsi 08:15:26.63394096 +45:35:32.011782 0.051 2018-04-09 3.3 140 , 206 , 233 , 340 , 379
2022dbl 12:20:45.078024 +49:33:04.40064 0.0284 2018-05-16 3.3 299 , 322 , 330 , 342 , 379
2018dyk 15:33:08.02008 +44:32:07.9188 0.03672 2018-05-31 2.0 123 , 134 , 143 , 157 , 203 , 368
2018cqh 02:33:46.93 -01:01:28.38 0.048000 2018-06-16 1.0 141 , 369 , 378
2018fyk 22:50:16.09153704 -44:51:53.50214628 0.059 2018-09-08 3.3 124 , 135 , 144 , 161 , 192 , 198 , 202 , 237 , 346 , 371 , 373 , 376 , 379
2018hco 01:07:33.63015048 +23:28:34.234662 0.088 2018-10-04 3.3 74 , 98 , 137 , 145 , 233 , 373 , 376 , 379
2018hyz 10:06:50.8848 +01:41:33.9144 0.0457266 2018-10-14 3.3 74 , 98 , 125 , 136 , 146 , 197 , 199 , 233 , 283 , 290 , 340 , 371 , 376 , 379 , 380
2018iih 17:28:03.92345736 +30:41:31.5128184 0.212 2018-11-09 3.3 147 , 205 , 233 , 339 , 376 , 379
2018jbv 13:10:45.55056 +08:34:04.2024 0.340000 2018-11-26 3.3 148 , 336 , 379
2018lni 04:09:37.65216 +73:53:41.658 0.138000 2018-11-28 3.3 173 , 233 , 340 , 379
2018lna 07:03:18.64896 +23:01:44.6988 0.091 2018-12-28 3.3 170 , 205 , 233 , 340 , 373 , 376 , 379
AT2019ahk 07:00:11.39064 -66:02:24.7704 0.0262 2019-01-29 3.3 158 , 169 , 375 , 376 , 379
2019cho 12:55:09.20976 +49:31:09.9264 0.193000 2019-02-12 3.3 174 , 233 , 379
2019bhf 15:09:15.97536 +16:14:22.524 0.1206 2019-02-12 3.3 172 , 205 , 233 , 376 , 379
2019azh 08:13:16.95037104 +22:38:53.8854 0.022 2019-02-22 3.3 151 , 163 , 168 , 171 , 204 , 233 , 237 , 241 , 284 , 288 , 292 , 339 , 346 , 371 , 372 , 376 , 379 , 386
2019cmw 18:48:39.4788 +51:00:48.7332 0.516700 2019-03-25 3.3 175 , 284 , 379
2019dsg 20:57:02.97 +14:12:15.86 0.0512 2019-04-09 3.3 53 , 164 , 176 , 185 , 233 , 238 , 245 , 287 , 339 , 346 , 371 , 373 , 376 , 379
2019ehz 14:09:41.90808 +55:29:27.78 0.074 2019-04-29 3.3 166 , 177 , 233 , 340 , 371 , 373 , 379
2019eve 11:28:49.66368 +15:40:22.3824 0.081300 2019-05-05 3.3 178 , 233 , 373 , 379
2019gte 13:06:39.424272 -01:32:54.41928 0.086000 2019-05-31 3.3 165 , 179 , 376
2019lwu 23:11:12.3054024 -01:00:10.8162072 0.117 2019-07-24 3.3 180 , 205 , 233 , 376 , 379
2019meg 18:45:16.19928 +44:26:18.9816 0.152000 2019-07-28 3.3 181 , 233 , 379
2019mha 16:16:27.78640272 +56:25:56.353224 0.148000 2019-07-30 3.3 182 , 233 , 379
2019qiz 04:46:37.88088 -10:13:34.878 0.0151 2019-09-19 3.3 74 , 98 , 167 , 184 , 200 , 224 , 233 , 340 , 376 , 379
2019teq 18:59:05.49192 +47:31:05.6136 0.0874 2019-10-20 3.3 186 , 208 , 336 , 371 , 373 , 379
2019vcb 12:38:56.38219704 +33:09:57.2151492 0.088 2019-11-15 3.3 187 , 204 , 336 , 371 , 376 , 379
2020pj 15:31:34.96759584 +33:05:41.6801292 0.068 2020-01-02 3.3 208 , 212 , 336 , 376 , 379
2020ddv 09:58:33.41648976 +46:54:40.4297676 0.16 2020-02-21 3.3 214 , 336 , 340 , 371 , 379
2020ksf 21:35:27.28176 -18:16:35.364 0.0942 2020-04-21 3.3 190 , 215 , 370 , 371 , 379
2020ocn 13:53:53.80106952 +53:59:49.7309352 0.0705 2020-04-29 3.3 188 , 189 , 213 , 220 , 232 , 235 , 336 , 340 , 346 , 371 , 379
2020mbq 15:40:15.26136 +25:00:04.8276 0.093000 2020-06-09 3.3 216 , 336 , 379
2020mot 00:31:13.58808 +85:00:31.9644 0.07 2020-06-14 3.3 207 , 217 , 336 , 339 , 379
2020neh 15:21:20.08237032 +14:04:10.5269304 0.062 2020-06-19 3.3 204 , 218 , 293 , 339 , 341 , 373 , 376 , 379
SRGE J135514.8+311605 13:55:15.01896 +31:16:05.2356 0.1989 2020-06-23 0.0 243 , 346
2020nov 16:58:12.97008 +02:07:03.0396 0.084 2020-06-27 3.3 211 , 219 , 341
SRGE J013204.6+122236 01:32:04.482 +12:22:35.6232 0.132 2020-07-08 0.0 243 , 346
2020opy 15:56:25.72872 +23:22:20.676 0.159 2020-07-08 3.3 221 , 336 , 341 , 343 , 379
SRGE J153503.4+455056 15:35:03.16128 +45:50:54.9528 0.2314 2020-07-13 0.0 243 , 346
2020qhs 02:17:53.95008 -09:36:50.8932 0.345000 2020-07-26 3.3 222 , 336 , 340 , 379
2020riz 02:10:30.75504 +09:04:26.7744 0.434000 2020-07-31 3.3 223 , 336 , 379
SRGE J163831.7+534020 16:38:32.01624 +53:40:22.5516 0.581 2020-08-01 0.0 243 , 346
SRGE J163030.2+470125 16:30:30.25248 +47:01:25.428 0.294 2020-08-06 0.0 243 , 346
SRGE J021939.9+361819 02:19:39.90336 +36:18:18.1944 0.3879 2020-08-07 0.0 243 , 346
SRGE J161001.2+330121 16:10:01.4208 +33:01:20.6976 0.1309 2020-08-14 0.0 243 , 346
SRGE J171423.6+085236 17:14:23.61432 +08:52:36.9048 0.036 2020-09-18 0.0 243 , 346
2020vdq 10:08:53.4409944 +42:43:00.22242 0.044 2020-10-04 3.3 225 , 339 , 353 , 354 , 357 , 366 , 379 , 386
2020wey 09:05:25.90992 +61:48:09.1368 0.02741 2020-10-08 3.3 209 , 226 , 336 , 339 , 379
AT2020vwl 15:30:37.8 +15:30:37.8 0.035000 2020-10-10 3.3 345 , 376
SRGE J071310.6+725627 07:13:10.5204 +72:56:26.7036 0.104 2020-10-11 0.0 243 , 346
SRGE J095928.6+643023 09:59:28.4784 +64:30:21.7908 0.454 2020-10-28 0.0 243 , 346
SRGE J091747.6+524821 09:17:47.39808 +52:48:20.286 0.1876 2020-10-28 0.0 243 , 346
2020ysg 11:25:26.0256 +27:26:26.1564 0.277000 2020-10-28 3.3 227 , 336 , 379
SRGE J133053.3+734824 13:30:53.02032 +73:48:24.2604 0.1505 2020-11-02 0.0 243 , 346
2020yue 11:00:00.336768 +21:06:45.95292 0.204 2020-11-03 3.3 228 , 304 , 339 , 358 , 379
2020zso 22:22:17.12592 -07:15:58.9392 0.0565 2020-11-12 3.3 210 , 229 , 282 , 336 , 340 , 376 , 379
SRGE J144738.4+671821 14:47:39.06504 +67:18:18.3384 0.125 2020-11-19 0.0 243 , 346
eRASSt J234403-352640 23:44:02.9 -35:26:39.4 0.100000 2020-11-28 3.1 335 , 381
2020abri 13:29:17.275152 +19:40:15.5082 0.178 2020-12-01 3.3 230 , 339 , 358 , 379
2020acka 15:55:01.947912 +16:18:16.33932 0.338 2020-12-11 3.3 231 , 247 , 339 , 358 , 376 , 379
2021lo 03:04:26.33856 +04:44:44.43684 0.152 2021-01-04 3.3 263 , 303 , 376 , 379
2021jsg 11:08:36.714696 +30:45:40.58424 0.126 2021-01-07 3.3 251 , 271 , 376 , 379
2021ack 13:53:09.98844 +09:44:25.34712 0.133 2021-01-12 3.3 249 , 264 , 376 , 379
2021axu 11:46:36.368448 +30:05:07.63584 0.19 2021-01-17 3.3 250 , 265 , 339 , 358 , 376 , 379
2021blz 04:32:31.2516 -32:25:50.367 0.045000 2021-01-31 3.3 248 , 266 , 376
2021crk 11:45:06.907392 +18:32:25.32408 0.155 2021-02-07 3.3 267 , 339 , 358 , 379
2021ehb 03:07:47.826648 +40:18:40.44816 0.017 2021-03-01 3.3 256 , 268 , 289 , 313 , 339 , 341 , 371 , 376 , 379
2021gje 16:50:07.324512 +34:49:07.70988 0.358 2021-03-18 3.3 253 , 254 , 269 , 376 , 379
2021jjm 14:39:30.666 -27:51:30.65688 0.153 2021-04-08 3.3 252 , 270 , 339 , 376 , 379
2021qxv 15:18:59.3 -03:11:45 0.183 2021-05-10 3.3 257 , 274 , 341 , 376
2021nwa 15:53:51.26957064 +55:35:19.82688 0.047 2021-05-11 3.3 255 , 273 , 339 , 358 , 379
2021mhg 00:19:42.893496 +29:19:00.71436 0.073 2021-05-11 3.3 258 , 272 , 339 , 376 , 379
2021sdu 01:11:23.84856 +50:34:29.70912 0.059 2021-07-05 3.3 259 , 275 , 339 , 358 , 379
2021uvz 17:32:50.521776 +33:01:58.1862 0.176 2021-07-12 3.3 278 , 307 , 376 , 379
2021uqv 00:32:39.876888 +22:32:56.15448 0.106 2021-07-29 3.3 261 , 276 , 339 , 358 , 376 , 379
2021utq 15:18:29.12088 +73:21:31.4136 0.127 2021-08-01 3.3 260 , 277 , 339 , 358 , 379
2021yzv 07:01:06.603816 +40:49:30.55908 0.286 2021-08-16 3.3 280 , 297 , 339 , 358 , 371 , 376 , 379
2021yte 06:55:04.733328 +12:38:03.17652 0.053 2021-09-12 3.3 262 , 279 , 339 , 376 , 379
2022rz 15:37:57.603648 +56:16:17.82876 0.107 2022-01-13 3.3 306 , 317 , 379
2022adm 14:27:51.43992 +28:10:29.0316 0.061000 2022-01-21 3.3 298 , 318 , 376
2022arb 10:24:10.33129176 -00:49:22.79202954 0.062000 2022-01-28 3.3 305 , 320 , 376
2022bdw 08:25:10.374984 +18:34:57.47304 0.0378 2022-01-31 3.3 300 , 321 , 342 , 376 , 379
2022cmc 13:34:43.207 +33:13:00.54 1.19325 2022-02-11 3.3 294 , 323 , 344
2022czy 12:22:01.12992 +16:59:44.0196 0.109000 2022-02-22 3.3 301 , 302 , 310 , 324 , 376
2022dyt 10:00:08.033544 +26:27:38.43828 0.072 2022-02-26 3.3 308 , 326 , 342 , 376 , 379
2022dsb 15:42:21.751824 -22:40:14.08764 0.023 2022-03-01 3.3 301 , 325 , 376 , 379
2022exr 17:29:50.510976 +25:50:31.79148 0.0955 2022-03-20 3.3 311 , 327 , 376 , 379
2022gri 07:18:20.794272 +33:59:41.47116 0.028 2022-04-03 3.3 295 , 309 , 328 , 379
2022hvp 09:54:45.24528 +55:26:25.4004 0.112 2022-04-19 3.3 296 , 312 , 329 , 379
2022wyq 09:48:06.55 +03:18:01.9 0.207 2022-04-26 1.0 74 , 98 , 114 , 152 , 334 , 346
2022lri 02:20:08.01192 -22:43:15.2472 0.03275 2022-06-01 3.3 315 , 331 , 377 , 379 , 384
2022pna 01:41:55.5504 -03:17:22.5204 0.095 2022-07-24 3.3 314 , 332 , 379
2022wtn 23:23:23.78448 +10:41:07.8108 0.0491 2022-10-02 3.3 316 , 333 , 379
2023clx 11:40:09.40296 +15:19:38.55 0.011070 2023-02-22 3.3 338 , 351 , 361 , 367 , 379
2023cvb 19:14:25.68072 +41:40:09.3036 0.071 2023-03-06 3.3 347 , 352 , 362 , 379
2023mfm 21:37:27.95784 -04:20:43.0152 0.087 2023-06-28 3.3 349 , 359 , 363 , 379
2023mhs 13:43:15.68112 +19:15:00.90324 0.0482 2023-06-30 3.3 348 , 355 , 356 , 364 , 379
2023rvb 19:21:12.549168 -22:38:39.60132 0.0815 2023-09-04 3.3 350 , 360 , 365 , 379
2024qab 18:14:06.192 +36:14:59.69472 0.1398 2024-06-15 3.3 379 , 383 , 385
AT2018dyb 16:10:58.87728 -60:55:24.3876 0.018 3.3 142 , 159 , 194 , 202 , 373 , 376 , 379
AT2022mim 16:45:54.69312 -23:27:06.9912 0.020350 3.2 368
J094608 09:46:08.49 +35:12:22.4 0.119 1.0 74 , 98 , 114 , 152 , 346
J233454 23:34:54.07 +14:57:12.8 0.107 1.0 74 , 98 , 114 , 152 , 346
SDSSJ1350 13:50:01.51 +29:16:09.71 0.0777 3.1 34 , 45 , 74 , 98 , 100 , 346
VAST J011148.1-025539 01:11:48 -02:55:48 1.0 374
VAST J015856.8-012404 01:58:57.6 -01:24:$0$ 1.0 374
VAST J093634.7-054755 09:36:33.6 -05:48:$0$ 1.0 374
VAST J104315.9+005059 10:43:16.8 +00:51:$0$ 1.0 374
VAST J144848.2+030235 14:48:48 +03:02:24 1.0 374
VAST J210626.2-020055 21:06:26.4 -02:01:12 1.0 374
[…]

### v2

@@ Appendix A Metadata and References for each TDE Candidate
ASASSN-15lh 22:02:15.49 -61:39:33.6 0.2326 2015-06-14 3.3 41 , 51 , 59 , 74 , 84 , 86 , 98 , 100 , 120 , 351 , 391
ASASSN-15oi 20:39:08.93 -30:45:22.6 0.0484 2015-08-14 3.3 5 , 60 , 74 , 83 , 98 , 100 , 105 , 198 , 201 , 244 , 351 , 382 , 387 , 391 , 401
OGLE16aaa 01:07:20.76144 -64:16:20.4456 0.1655 2015-12-17 3.3 5 , 74 , 98 , 100 , 107 , 108 , 351 , 391
Gaia 16aax 14:34:18 +12:49:37 0.248 2016-01-26 3.3 107 , 196 , 351
Gaia 16ajq 14:53:02 +42:21:28 0.28 2016-03-29 3.3 88 , 89 , 107 , 130 , 351
Gaia 16aka 14:48:08 +37:13:59 0.31 2016-03-31 3.3 88 , 107 , 351
iPTF16axa 17:03:34.36 +30:35:36.8 0.108 2016-05-29 3.3 5 , 74 , 98 , 101 , 111 , 351 , 384 , 391
AT2016ezh 01:58:05 -00:52:22 0.0804 2016-08-26 3.3 5 , 51 , 74 , 90 , 98 , 103 , 114 , 202 , 236 , 351 , 384
2016fnl 00:29:57.04 +32:53:37.5 0.0163 2016-08-26 3.3 5 , 74 , 75 , 87 , 91 , 98 , 100 , 104 , 111 , 239 , 351 , 384 , 387 , 391
OGLE17aaj 01:56:24.93 -71:04:15.7 0.116 2017-01-01 3.1 74 , 94 , 97 , 98 , 149 , 351
2017bcc 11:31:52.9875888 +29:59:44.8267578 0.133 2017-02-18 3.3 74 , 97 , 98 , 117 , 118 , 351 , 394
2019pqt 18:08:07.549 +13:15:05.4 2017-03-15 0.0 95 , 96 , 98 , 183 , 242 , 351
2017eqx 22:26:48.37042128 +17:08:52.39825116 0.1089 2017-06-01 3.3 74 , 98 , 119 , 160 , 202 , 351 , 384 , 391
2018gn 01:46:42.4356 +32:30:29.4048 0.037500 2018-01-14 3.3 133 , 138 , 378
XMMSL2 J140446.9-251135 14:04:47 -25:11:35 0.043 2018-02-15 1.0 122 , 351
2018zr 07:56:54.54888 +34:15:43.6284 0.071 2018-03-06 3.3 74 , 98 , 121 , 139 , 153 , 155 , 156 , 202 , 205 , 345 , 351 , 382 , 384 , 387 , 391
2018bsi 08:15:26.63394096 +45:35:32.011782 0.051 2018-04-09 3.3 140 , 206 , 233 , 345 , 391
2022dbl 12:20:45.078024 +49:33:04.40064 0.0284 2018-05-16 3.3 300 , 324 , 334 , 347 , 391
2018dyk 15:33:08.02008 +44:32:07.9188 0.03672 2018-05-31 2.0 123 , 134 , 143 , 157 , 203 , 378
2018cqh 02:33:46.93 -01:01:28.38 0.048000 2018-06-16 1.0 141 , 379 , 390
2018fyk 22:50:16.09153704 -44:51:53.50214628 0.059 2018-09-08 3.3 124 , 135 , 144 , 161 , 192 , 198 , 202 , 237 , 351 , 382 , 384 , 387 , 391
2018hco 01:07:33.63015048 +23:28:34.234662 0.088 2018-10-04 3.3 74 , 98 , 137 , 145 , 233 , 384 , 387 , 391
2018hyz 10:06:50.8848 +01:41:33.9144 0.0457266 2018-10-14 3.3 74 , 98 , 125 , 136 , 146 , 197 , 199 , 233 , 284 , 291 , 345 , 382 , 387 , 391 , 392
2018iih 17:28:03.92345736 +30:41:31.5128184 0.212 2018-11-09 3.3 147 , 205 , 233 , 344 , 387 , 391
2018jbv 13:10:45.55056 +08:34:04.2024 0.340000 2018-11-26 3.3 148 , 341 , 391
2018lni 04:09:37.65216 +73:53:41.658 0.138000 2018-11-28 3.3 173 , 233 , 345 , 391
2018lna 07:03:18.64896 +23:01:44.6988 0.091 2018-12-28 3.3 170 , 205 , 233 , 345 , 384 , 387 , 391
AT2019ahk 07:00:11.39064 -66:02:24.7704 0.0262 2019-01-29 3.3 158 , 169 , 386 , 387 , 391
2019cho 12:55:09.20976 +49:31:09.9264 0.193000 2019-02-12 3.3 174 , 233 , 391
2019bhf 15:09:15.97536 +16:14:22.524 0.1206 2019-02-12 3.3 172 , 205 , 233 , 387 , 391
2019azh 08:13:16.95037104 +22:38:53.8854 0.022 2019-02-22 3.3 151 , 163 , 168 , 171 , 204 , 233 , 237 , 241 , 285 , 289 , 293 , 344 , 351 , 382 , 383 , 387 , 391 , 400
2019cmw 18:48:39.4788 +51:00:48.7332 0.516700 2019-03-25 3.3 175 , 285 , 391
2019dsg 20:57:02.97 +14:12:15.86 0.0512 2019-04-09 3.3 53 , 164 , 176 , 185 , 233 , 238 , 245 , 288 , 344 , 351 , 382 , 384 , 387 , 391
2019ehz 14:09:41.90808 +55:29:27.78 0.074 2019-04-29 3.3 166 , 177 , 233 , 345 , 382 , 384 , 391
2019eve 11:28:49.66368 +15:40:22.3824 0.081300 2019-05-05 3.3 178 , 233 , 384 , 391
2019gte 13:06:39.424272 -01:32:54.41928 0.086000 2019-05-31 3.3 165 , 179 , 387
2019lwu 23:11:12.3054024 -01:00:10.8162072 0.117 2019-07-24 3.3 180 , 205 , 233 , 387 , 391
2019meg 18:45:16.19928 +44:26:18.9816 0.152000 2019-07-28 3.3 181 , 233 , 391
2019mha 16:16:27.78640272 +56:25:56.353224 0.148000 2019-07-30 3.3 182 , 233 , 391
2019qiz 04:46:37.88088 -10:13:34.878 0.0151 2019-09-19 3.3 74 , 98 , 167 , 184 , 200 , 224 , 233 , 345 , 387 , 391
2019teq 18:59:05.49192 +47:31:05.6136 0.0874 2019-10-20 3.3 186 , 208 , 341 , 382 , 384 , 391
2019vcb 12:38:56.38219704 +33:09:57.2151492 0.088 2019-11-15 3.3 187 , 204 , 341 , 382 , 387 , 391
2020pj 15:31:34.96759584 +33:05:41.6801292 0.068 2020-01-02 3.3 208 , 212 , 341 , 387 , 391
2020ddv 09:58:33.41648976 +46:54:40.4297676 0.16 2020-02-21 3.3 214 , 341 , 345 , 382 , 391
2020ksf 21:35:27.28176 -18:16:35.364 0.0942 2020-04-21 3.3 190 , 215 , 380 , 382 , 391
2020ocn 13:53:53.80106952 +53:59:49.7309352 0.0705 2020-04-29 3.3 188 , 189 , 213 , 220 , 232 , 235 , 341 , 345 , 351 , 382 , 391
2020mbq 15:40:15.26136 +25:00:04.8276 0.093000 2020-06-09 3.3 216 , 341 , 391
2020mot 00:31:13.58808 +85:00:31.9644 0.07 2020-06-14 3.3 207 , 217 , 341 , 344 , 391
2020neh 15:21:20.08237032 +14:04:10.5269304 0.062 2020-06-19 3.3 204 , 218 , 294 , 344 , 346 , 384 , 387 , 391
SRGE J135514.8+311605 13:55:15.01896 +31:16:05.2356 0.1989 2020-06-23 0.0 243 , 351
2020nov 16:58:12.97008 +02:07:03.0396 0.084 2020-06-27 3.3 211 , 219 , 346
SRGE J013204.6+122236 01:32:04.482 +12:22:35.6232 0.132 2020-07-08 0.0 243 , 351
2020opy 15:56:25.72872 +23:22:20.676 0.159 2020-07-08 3.3 221 , 341 , 346 , 348 , 391
SRGE J153503.4+455056 15:35:03.16128 +45:50:54.9528 0.2314 2020-07-13 0.0 243 , 351
2020qhs 02:17:53.95008 -09:36:50.8932 0.345000 2020-07-26 3.3 222 , 341 , 345 , 391
2020riz 02:10:30.75504 +09:04:26.7744 0.434000 2020-07-31 3.3 223 , 341 , 391
SRGE J163831.7+534020 16:38:32.01624 +53:40:22.5516 0.581 2020-08-01 0.0 243 , 351
SRGE J163030.2+470125 16:30:30.25248 +47:01:25.428 0.294 2020-08-06 0.0 243 , 351
SRGE J021939.9+361819 02:19:39.90336 +36:18:18.1944 0.3879 2020-08-07 0.0 243 , 351
SRGE J161001.2+330121 16:10:01.4208 +33:01:20.6976 0.1309 2020-08-14 0.0 243 , 351
SRGE J171423.6+085236 17:14:23.61432 +08:52:36.9048 0.036 2020-09-18 0.0 243 , 351
2020vdq 10:08:53.4409944 +42:43:00.22242 0.044 2020-10-04 3.3 225 , 344 , 358 , 359 , 363 , 391 , 400
2020wey 09:05:25.90992 +61:48:09.1368 0.02741 2020-10-08 3.3 209 , 226 , 341 , 344 , 391
AT2020vwl 15:30:37.8 +26:58:56.89 0.035000 2020-10-10 3.3 350 , 387
SRGE J071310.6+725627 07:13:10.5204 +72:56:26.7036 0.104 2020-10-11 0.0 243 , 351
SRGE J095928.6+643023 09:59:28.4784 +64:30:21.7908 0.454 2020-10-28 0.0 243 , 351
SRGE J091747.6+524821 09:17:47.39808 +52:48:20.286 0.1876 2020-10-28 0.0 243 , 351
2020ysg 11:25:26.0256 +27:26:26.1564 0.277000 2020-10-28 3.3 227 , 341 , 391
SRGE J133053.3+734824 13:30:53.02032 +73:48:24.2604 0.1505 2020-11-02 0.0 243 , 351
2020yue 11:00:00.336768 +21:06:45.95292 0.204 2020-11-03 3.3 228 , 305 , 344 , 364 , 391
2020zso 22:22:17.12592 -07:15:58.9392 0.0565 2020-11-12 3.3 210 , 229 , 283 , 341 , 345 , 387 , 391
SRGE J144738.4+671821 14:47:39.06504 +67:18:18.3384 0.125 2020-11-19 0.0 243 , 351
eRASSt J234403-352640 23:44:02.9 -35:26:39.4 0.100000 2020-11-28 3.1 340 , 393
2020abri 13:29:17.275152 +19:40:15.5082 0.178 2020-12-01 3.3 230 , 344 , 364 , 391
2020acka 15:55:01.947912 +16:18:16.33932 0.338 2020-12-11 3.3 231 , 247 , 344 , 364 , 387 , 391
2021lo 03:04:26.33856 +04:44:44.43684 0.152 2021-01-04 3.3 263 , 304 , 387 , 391
2021jsg 11:08:36.714696 +30:45:40.58424 0.126 2021-01-07 3.3 251 , 271 , 387 , 391
2021ack 13:53:09.98844 +09:44:25.34712 0.133 2021-01-12 3.3 249 , 264 , 387 , 391
2021axu 11:46:36.368448 +30:05:07.63584 0.19 2021-01-17 3.3 250 , 265 , 344 , 364 , 387 , 391
2021blz 04:32:31.2516 -32:25:50.367 0.045000 2021-01-31 3.3 248 , 266 , 387
2021crk 11:45:06.907392 +18:32:25.32408 0.155 2021-02-07 3.3 267 , 344 , 364 , 391
2021ehb 03:07:47.826648 +40:18:40.44816 0.017 2021-03-01 3.3 256 , 268 , 290 , 315 , 344 , 346 , 382 , 387 , 391
2021gje 16:50:07.324512 +34:49:07.70988 0.358 2021-03-18 3.3 253 , 254 , 269 , 387 , 391
2021jjm 14:39:30.666 -27:51:30.65688 0.153 2021-04-08 3.3 252 , 270 , 344 , 387 , 391
2021qxv 15:18:59.3 -03:11:45 0.183 2021-05-10 3.3 257 , 275 , 346 , 387
2021nwa 15:53:51.26957064 +55:35:19.82688 0.047 2021-05-11 3.3 255 , 273 , 344 , 364 , 391
2021mhg 00:19:42.893496 +29:19:00.71436 0.073 2021-05-11 3.3 258 , 272 , 344 , 387 , 391
2021qth 20:11:38.92584 -21:09:36.6444 0.080500 2021-06-17 3.0 274 , 364
2021sdu 01:11:23.84856 +50:34:29.70912 0.059 2021-07-05 3.3 259 , 276 , 344 , 364 , 391
2021uvz 17:32:50.521776 +33:01:58.1862 0.176 2021-07-12 3.3 279 , 308 , 387 , 391
2021uqv 00:32:39.876888 +22:32:56.15448 0.106 2021-07-29 3.3 261 , 277 , 344 , 364 , 387 , 391
2021utq 15:18:29.12088 +73:21:31.4136 0.127 2021-08-01 3.3 260 , 278 , 344 , 364 , 391
2021yzv 07:01:06.603816 +40:49:30.55908 0.286 2021-08-16 3.3 281 , 298 , 344 , 364 , 382 , 387 , 391
2021yte 06:55:04.733328 +12:38:03.17652 0.053 2021-09-12 3.3 262 , 280 , 344 , 387 , 391
2022rz 15:37:57.603648 +56:16:17.82876 0.107 2022-01-13 3.3 307 , 319 , 391
2022adm 14:27:51.43992 +28:10:29.0316 0.061000 2022-01-21 3.3 299 , 320 , 387
2022arb 10:24:10.33129176 -00:49:22.79202954 0.062000 2022-01-28 3.3 306 , 322 , 387
2022bdw 08:25:10.374984 +18:34:57.47304 0.0378 2022-01-31 3.3 301 , 323 , 347 , 387 , 391
2022cmc 13:34:43.207 +33:13:00.54 1.19325 2022-02-11 3.3 295 , 325 , 349
2022csn 07:28:54.93 +26:53:49.3116 0.148000 2022-02-19 3.0 326
2022czy 12:22:01.12992 +16:59:44.0196 0.109000 2022-02-22 3.3 302 , 303 , 311 , 327 , 387
2022dyt 10:00:08.033544 +26:27:38.43828 0.072 2022-02-26 3.3 309 , 329 , 347 , 387 , 391
2022dsb 15:42:21.751824 -22:40:14.08764 0.023 2022-03-01 3.3 302 , 328 , 387 , 391
2022exr 17:29:50.510976 +25:50:31.79148 0.0955 2022-03-20 3.3 312 , 330 , 387 , 391
2022fpx 15:31:03.70008 +53:24:19.26 0.073000 2022-03-31 3.0 314 , 331
2022gri 07:18:20.794272 +33:59:41.47116 0.028 2022-04-03 3.3 296 , 310 , 332 , 391
2022hvp 09:54:45.24528 +55:26:25.4004 0.112 2022-04-19 3.3 297 , 313 , 333 , 391
2022wyq 09:48:06.55 +03:18:01.9 0.207 2022-04-26 1.0 74 , 98 , 114 , 152 , 339 , 351
2022lri 02:20:08.01192 -22:43:15.2472 0.03275 2022-06-01 3.3 317 , 335 , 388 , 391 , 396
2022pna 01:41:55.5504 -03:17:22.5204 0.095 2022-07-24 3.3 316 , 336 , 391
2022upj 00:23:56.84592 -14:25:23.2212 0.054000 2022-08-31 3.0 337 , 389
2022wtn 23:23:23.78448 +10:41:07.8108 0.0491 2022-10-02 3.3 318 , 338 , 391
2023clx 11:40:09.40296 +15:19:38.55 0.011070 2023-02-22 3.3 343 , 356 , 369 , 377 , 391
2023cvb 19:14:25.68072 +41:40:09.3036 0.071 2023-03-06 3.3 352 , 357 , 370 , 391
2023kvy 23:39:47.96376 +08:09:53.6256 0.160000 2023-05-28 3.0 366 , 371
2023lli 22:57:39.46992 +40:32:40.02 0.036000 2023-06-23 3.0 360 , 372 , 381
2023mfm 21:37:27.95784 -04:20:43.0152 0.087 2023-06-28 3.3 354 , 365 , 373 , 391
2023mhs 13:43:15.68112 +19:15:00.90324 0.0482 2023-06-30 3.3 353 , 361 , 362 , 374 , 391
2023rvb 19:21:12.549168 -22:38:39.60132 0.0815 2023-09-04 3.3 355 , 368 , 375 , 391
2023ugy 21:40:44.02008 +21:00:58.5684 0.106000 2023-10-06 3.0 367 , 376
2024qab 18:14:06.192 +36:14:59.69472 0.1398 2024-06-15 3.3 391 , 395 , 398
2024tvd 17:10:42.58488 +28:50:15.108 0.044938 2024-08-25 3.0 397 , 399 , 402 , 403
AT2018dyb 16:10:58.87728 -60:55:24.3876 0.018 3.3 142 , 159 , 194 , 202 , 384 , 387 , 391
AT2022mim 16:45:54.69312 -23:27:06.9912 0.020350 3.2 378
J094608 09:46:08.49 +35:12:22.4 0.119 1.0 74 , 98 , 114 , 152 , 351
J233454 23:34:54.07 +14:57:12.8 0.107 1.0 74 , 98 , 114 , 152 , 351
SDSSJ1350 13:50:01.51 +29:16:09.71 0.0777 3.1 34 , 45 , 74 , 98 , 100 , 351
VAST J011148.1-025539 01:11:48 -02:55:48 1.0 385
VAST J015856.8-012404 01:58:57.6 -01:24:$0$ 1.0 385
VAST J093634.7-054755 09:36:33.6 -05:48:$0$ 1.0 385
VAST J104315.9+005059 10:43:16.8 +00:51:$0$ 1.0 385
VAST J144848.2+030235 14:48:48 +03:02:24 1.0 385
[…]